OpenStudy (anonymous):

f(x)=-2x-7, g(x)=-4x3, find f(g(x))

OpenStudy (luigi0210):

f(g(x)) is another way of saying plug in the equation of g(x) in for x in f(x)

OpenStudy (anonymous):

is that 4x^3

OpenStudy (luigi0210):

wait, what is the g(x) function?

OpenStudy (luigi0210):

If it is a 4x^3 then it would look like this. \[f(4x^3)=-2(4x^3)-7\] Now multiply and simplify \[f(4x^3)=-8x^3-7\]

OpenStudy (luigi0210):

Hope that helps

OpenStudy (anonymous):

thanks